OpenStudy (anonymous):

Can someone help me please? Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1) A. y + 3 = -5(x - 2) B. y - 2 = -5(x + 3) C. y + 3 = -1/5(x-2) D. y - 2 =-1/5(x+3)

OpenStudy (anonymous):

@gorv , can you help you seem to bee the only person willing to help me?

OpenStudy (paxpolaris):

Find the slope

OpenStudy (paxpolaris):

\[m={rise \over run}={y_2-y_1 \over x_2=x_1}\]

OpenStudy (gorv):

point slope form is y-y1=[(y2-y1)/(x2-x1)](x-x1)

OpenStudy (anonymous):

so it would be d?

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

ok than you!

OpenStudy (anonymous):

Can i get help on one more @gorv

OpenStudy (anonymous):

Given the equation y - 3 = (x + 6) in point-slope form, identify the equation of the same line in standard form. A. x - 2y = -12 B. y = 1/2x C. y = 1/2 x + 6 D. y = 1/2 x + 9

OpenStudy (gorv):

add 3 on both side

OpenStudy (anonymous):

d?

OpenStudy (gorv):

i think u typed the Q wrong

OpenStudy (anonymous):

oh i am sorry the question should be y-3=1/2(x+6)?